# Local Thermal Operations and Classical Communication

### v2

@@ 5.2 Bithermal tensors approaching zero temperature
The extremal thermal and bithermal tensors are the ones with only 00 and $1$ entries, restricted only by the abovestated condition.
@@
5.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
Although the general construction of extremal bithermal tensors seems unattainable, based on the results introduced in [62] we may construct a family of extreme bithermal tensors, which exist for every Hamiltonian when the temperature is low enough. Hereafter, for the sake of clarity, we adopt the notation $g_{i,j}=e^{-\beta(E_{i}-E_{j})}$ [62]. In dimension $2$ our example is just an extremal tensor (27)
$T_{ijk}|_{d=2}=(T_{i,0,k}|T_{i,1,k})_{d=2}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0\\ 0&1\\ \end{matrix}\right)\leavevmode\nobreak\ ,$ (35)
@@ 5.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
$T_{ijk}|_{d=3}$ $=(T_{i,0,k}|T_{i,1,k}|T_{i,2,k})|_{d=3}=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&1&1\\ g_{1,0}&0&0\\ g_{2,0}&0&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1-g_{2,1}&1\\ 0&g_{2,1}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right)=$ (36)
$=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&\vline&\begin{matrix}1&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ g_{2,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=2}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\vline&T_{i,1,k}|_{d=2}\end{matrix}\right)$
@@
$=\left(\begin{matrix}1-\sum_{i=1}^{n}g_{i,0}&\vline&\begin{matrix}1&\cdots&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ \vdots\\ g_{d,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=n-1}\end{matrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \right|\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,n-1,k}|_{d=n-1}\end{matrix}\right)\leavevmode\nobreak\ .$
@@
$\sum_{i=1}^{n}g_{i,0}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=2}^{n}g_{i,1}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \cdots,\leavevmode\nobreak\ g_{n-1,n-2}+g_{n,n-2}\leq 1\leavevmode\nobreak\ .$ (38)
@@
Figure 7: Graphical representation of extremal bithermal tensor (37) in dimension $d=4$ as a cube and its separation into consecutive slices. Grey cells correspond to value $1$, red-to-blue cells to other nonzero values, and empty cells to 00.
5.4 Majorization-like restrictions on outputs of (bi)thermal tensors
Having discussed the general structure of the set of thermal and bithermal tensors, including methods for obtaining the extreme points, we shift our attention to the images of pairs of states under LTOCC and $\mathcal{S}$LTOCC. We start our discussion by considering local results. In other words, we will consider what Bob can obtain in his subsystem after reciving information from Alice, excluding the correlations between subsystems. Thus, we will focus on one (bi)thermal tensor, rather than the entire scheme. We stress that although the entire sub-section is presented from the perspective of thermal operations, the presented results hold also for (tri)stochastic tensors.
@@ 5.4 Majorization-like restrictions on outputs of (bi)thermal tensors
The tensor $T$ is a [bi-]thermal tensor, $\forall_{q}$ $T({\mathbf{\gamma}},\mathbf{q})={\mathbf{\gamma}}$ [and $T({\mathbf{q},\mathbf{\gamma}})={\mathbf{\gamma}}$], if and only if for any input distributions $p,q$ the output is a probability distribution satisfying thermomajorization relations $\mathbf{p}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$ [and $\mathbf{q}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$].
@@
Let $T$ be a thermal tensor then for any distribution $\mathbf{q}$, $T(\cdot\leavevmode\nobreak\ ,\mathbf{q})$ is a thermal operation, thus the implication in one direction follows.
@@
On the other hand, if there is no convex combination of majorization curves that majorizes $\mathbf{r}$, there is also no combination of extremal tensors which would lead $\mathbf{p},\mathbf{q}$ into state distribution majorizing $\mathbf{r}$ thus there is no desired tensor. ∎
@@
Let $l$ be a curve obtained by, at each point, taking a maximum over majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$, with $T^{(i)}$ being once again extremal bithermal tensors. If $l$ does not majorize the curve corresponding to $\mathbf{r}$, then there exists no bithermal tensor $T$, such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$.
@@
$=(1-w_{p})(1-w_{q})T(\tilde{\mathbf{p}},\tilde{\mathbf{q}})+\left[1-(1-w_{p})(1-w_{q})\right]{\gamma}$
@@
Thus we effectively boosted the requirements proceeding from thermomajorization relations, leveraging the fact, that the allowed operations are bithermal.
6 Examples and properties of LTOCC protocols
In this section we will provide two examples of application of LTOCC and their extensions. In the first subsection we will consider the problem of approximating basic nonlocal logic gates – CNOT and SWAP – using LTOCC with memory restricted to each round of the protocol. Then, we will consider the case of parallel LTOCC and possibility of generating correlations between the subsystems.
6.1 Approximating logic gates using LTOCC
In this subsection, we provide examples of LTOCC maps without memory, which approximate standard logical operations on bits: CNOT and SWAP (between subsystems). Since both of those operations are nonlocal, they cannot be constructed from just local thermal operations. On the other hand, both of those operations have a straightforward construction using LOCC, which we present.
$\delta(CNOT,CNOT^{(\beta)})=g_{B}$
which, in the limit of large temperatures $g_{A},g_{B}\to 1$, both coincide with standard $SWAP$
$SWAP=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix}\leavevmode\nobreak\ ,$
and the statistical distance between them is given by
$\delta(SWAP,SWAP_{A}^{(\beta)})=3-g_{A}-2g_{B}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \delta(SWAP,SWAP_{B}^{(\beta)})=3-2g_{A}-g_{B}.$
6.2 Generating correlations from product input states
Finally, we return to the general discussion of one-round and two-round LTOCC between two parties to demonstrate its ability to generate correlations.
@@ 6.2 Generating correlations from product input states
Let $T_{jkl}$ be a thermal tensor, $\sum_{l}T_{jkl}\gamma_{l}=\gamma_{j}$, having an identity [or permutation matrix in infinite temperature case] as at least one layer $T_{jkl_{0}}=\delta_{j,k}$ or $T_{jkl_{0}}=\delta_{j,\sigma(k)}$ for some permutation $\sigma$]. Then the pair of states $\mathbf{p},\mathbf{q}$, second of which is a sharp state $q_{l}=\delta_{l,l_{0}}$ can be mapped into the maximally correlated state in one-round LTOCC.
@@
Let us take $\Lambda_{ik}=\delta_{i,k}$ (or $\Lambda_{ik}=\delta_{i,\sigma(k)}$ in infinite temperature) and define $M_{ij,kl}=\Lambda_{ik}T_{jkl}$, then the output distribution $\pi_{ij}$ is given by
$\pi_{ij}=\sum_{kl}M_{ij,kl}p_{k}q_{l}=\sum_{kl}\Lambda_{ik}T_{jkl}p_{k}\delta_{l,l_{0}}=\sum_{k}\delta_{i,k}\delta_{j,k}p_{k}=\delta_{i,j}p_{i}\leavevmode\nobreak\ ,$
@@
$H(A|B)=H(B|A)=-\sum_{ij}\pi_{ij}\ln\left(\frac{\pi_{ij}}{p_{i}}\right)=-\sum_{ij}\delta_{i,j}p_{i}\ln\left(\frac{p_{i}}{p_{i}}\right)=-\sum_{i}p_{i}\ln(1)=0.$ (43)
Notice that the requirement for thermal tensor demanded above can be guaranteed not only in zero or infinite temperature but is satisfied, for example, by construction (37). Additional discussion and examples of correlating properties of LTOCC in finite temperature are presented in the C.
Suppose we allow ourselves to use one round of parallel LTOCC, especially $\mathcal{S}$LTOCC, which is a subset of two rounds of LTOCC with memory. In that case the above observation can be extended to encompass non-sharp states, and even Gibbs state.
@@
Consider a $\mathcal{S}LTOCC$ operation consisting of two copies of the same extremal (bi)thermal tensor $M_{ij,kl}=T_{ikl}T_{jkl}$, which entries are either 00 or $1$. Then for any product input state $\mathbf{p}\otimes\mathbf{q}$ the output state $\mathbf{\pi}=M(\mathbf{p}\otimes\mathbf{q})$ is maximally correlated.
@@
Let us denote $T(\mathbf{p},\mathbf{q})=\mathbf{r}$, then, since in each column ($\forall_{j,k}$) of $T_{ijk}$ there is only a single entry equal to $1$ and the rest of them are 00, we get
$\pi_{ij}=\sum_{kl}M_{ij,lk}p_{k}q_{l}=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{kl}\delta_{i,j}T_{ijk}p_{k}q_{l}=\delta_{ij}r_{i}\leavevmode\nobreak\ .$ (44)
@@
7 Summary and Outlook

### v4

@@ V.2 Bithermal tensors approaching zero temperature
The extremal thermal and bithermal tensors are the ones with only 0 and $1$ entries, restricted only by the abovestated condition.
@@
V.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
Although the general construction of extremal bithermal tensors seems unattainable, based on the results introduced in [67] we may construct a particular family of extreme bithermal tensors, which exist for every Hamiltonian when the temperature is low enough. Hereafter, for the sake of clarity, we adopt the notation $g_{i,j}=e^{-\beta(E_{i}-E_{j})}$ [67]. In dimension $2$ our example is just an extremal tensor (27)
$T_{ijk}|_{d=2}=(T_{i,0,k}|T_{i,1,k})_{d=2}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0\\ 0&1\\ \end{matrix}\right)\penalty 10000\ ,$ (34)
@@ V.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
$T_{ijk}|_{d=3}$ $=(T_{i,0,k}|T_{i,1,k}|T_{i,2,k})|_{d=3}$ (35)
$=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&1&1\\ g_{1,0}&0&0\\ g_{2,0}&0&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1-g_{2,1}&1\\ 0&g_{2,1}&0\\ \end{matrix}\right.\left|\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{matrix}\right)=\left(\begin{matrix}1-g_{1,0}-g_{2,0}&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}1&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ g_{2,0}\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&\mbox{\Large 0}\end{matrix}\right.\penalty 10000\ \left|\penalty 10000\ \penalty 10000\ \begin{matrix}1&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&T_{i,0,k}|_{d=2}\end{matrix}\right.\penalty 10000\ \left|\penalty 10000\ \penalty 10000\ \begin{matrix}1&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}0&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ 0\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&T_{i,1,k}|_{d=2}\end{matrix}\right)$
@@
$=\left(\begin{matrix}1-\sum_{i=1}^{n}g_{i,0}&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}1&\cdots&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ \vdots\\ g_{d,0}\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&\mbox{\Large 0}\end{matrix}\right.\penalty 10000\ \left|\penalty 10000\ \penalty 10000\ \begin{matrix}1&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&T_{i,0,k}|_{d=n-1}\end{matrix}\penalty 10000\ \penalty 10000\ \right|\penalty 10000\ \penalty 10000\ \penalty 10000\ \cdots\penalty 10000\ \penalty 10000\ \penalty 10000\ \left|\penalty 10000\ \penalty 10000\ \begin{matrix}1&\hskip-5.0pt\vline\hskip-5.0pt&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\hskip-5.0pt\vline\hskip-5.0pt&T_{i,n-1,k}|_{d=n-1}\end{matrix}\right)\penalty 10000\ .$
@@
$\sum_{i=1}^{n}g_{i,0}\leq 1\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \sum_{i=2}^{n}g_{i,1}\leq 1\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \cdots,\penalty 10000\ g_{n-1,n-2}+g_{n,n-2}\leq 1\penalty 10000\ .$ (37)
@@
Figure 7: Graphical representation of extremal bithermal tensor (36) in dimension $d=4$ as a cube and its separation into consecutive slices. Grey cells correspond to value 11, red-to-blue cells to other nonzero values, and empty cells to 0.
V.4 Majorization-like restrictions on outputs of (bi)thermal tensors
Having discussed the general structure of the set of thermal and bithermal tensors, including methods for obtaining the extreme points, we shift our attention to the images of pairs of states under LTOCC and $\mathcal{S}$LTOCC. We start our discussion by considering local results. In other words, we will consider what Bob can obtain in his subsystem after reciving information from Alice, excluding the correlations between subsystems. Thus, we will focus on one (bi)thermal tensor, rather than the entire scheme. We stress that although the entire sub-section is presented from the perspective of thermal operations, the presented results hold also in the $\beta\to 0$ limit, corresponding to (tri)stochastic tensors.
@@ V.4 Majorization-like restrictions on outputs of (bi)thermal tensors
The tensor $T$ is a [bi-]thermal tensor, $\forall_{q}$ $T({\mathbf{\gamma}},\mathbf{q})={\mathbf{\gamma}}$ [and $T({\mathbf{q},\mathbf{\gamma}})={\mathbf{\gamma}}$], if and only if for any input distributions $\mathbf{p},\mathbf{q}$ the output is a probability distribution satisfying thermomajorization relations $\mathbf{p}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$ [and $\mathbf{q}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$].
@@
Let $T$ be a thermal tensor then for any distribution $\mathbf{q}$, $T(\cdot\penalty 10000\ ,\mathbf{q})$ is a thermal operation, thus the implication in one direction follows.
@@
On the other hand, if there is no convex combination of majorization curves that majorizes $\mathbf{r}$, there is also no combination of extremal tensors which would lead $\mathbf{p},\mathbf{q}$ into state distribution majorizing $\mathbf{r}$ - thus there is no desired tensor. ∎
@@
Let $l$ be a curve obtained by taking a maximum over majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$ at each point, with $T^{(i)}$ being extremal bithermal tensors. If $l$ does not majorize the curve corresponding to $\mathbf{r}$, then there exists no bithermal tensor $T$, such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$.
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ =(1-w_{p})(1-w_{q})T(\tilde{\mathbf{p}},\tilde{\mathbf{q}})+\left[1-(1-w_{p})(1-w_{q})\right]{\gamma}$
@@
Thus we effectively improve the requirements proceeding from thermomajorization relations, leveraging the fact, that the allowed operations are bithermal.
VI Nonlocal correlations under LTOCC
In the following two sections, we ask a natural questions concerning relation between LTOCC and nonlocality. First, we demonstrate that starting from initially uncorrelated energy-incoherent systems in product state one may achieve significant level of correlations between the systems, in particular demosntrating the power of LTOCC with memory. Next, we move to present two aspects of LTOCC in the CHSH scenario as an example of test of Bell nonlocality. We show that on one hand it is generically impossible to detect breaking of CHSH inequality in a single-copy scenario due to restrictions imposed by local thermal operations; on the other hand, when multiple copies are present, the upper bound achievable by LTOCC is below Tsirelson bound, and tends to it in the limit of infinite copies, leading to possiblity of detecting athermal resource usage in finite-copy setting.
VI.1 Generating correlations from product input states
First, we will focus on the ability of one-round and two-round LTOCC between two parties to demonstrate its ability to generate nonlocal correlations when acting on product energy-incoherent states.
@@ VI.1 Generating correlations from product input states
Let $T_{jkl}$ be a thermal tensor, $\sum_{l}T_{jkl}\gamma_{l}=\gamma_{j}$, having an identity [or permutation matrix in infinite temperature case] as at least one layer $T_{jkl_{0}}=\delta_{j,k}$ [or $T_{jkl_{0}}=\delta_{j,\sigma(k)}$ for some permutation $\sigma$]. Then the pair of states $\mathbf{p},\mathbf{q}$, second of which is a sharp state $q_{l}=\delta_{l,l_{0}}$ can be mapped into the maximally correlated state in one-round LTOCC.
@@
Let us take $\Lambda_{ik}=\delta_{i,k}$ (or $\Lambda_{ik}=\delta_{i,\sigma(k)}$ in infinite temperature) and consider one-round LTOCC $M_{ij,kl}=\Lambda_{ik}T_{jkl}$ acting on uncorelaterd distribution $\mathbf{p}\otimes\mathbf{q}$. Then the output distribution $\pi_{ij}$ is given by
$\pi_{ij}$ $=\sum_{kl}M_{ij,kl}p_{k}q_{l}=\sum_{kl}\Lambda_{ik}T_{jkl}p_{k}\delta_{l,l_{0}}=\sum_{k}\delta_{i,k}\delta_{j,k}p_{k}$
$=\delta_{i,j}p_{i}\penalty 10000\ ,\$
@@
$H(A|B)=H(B|A)=-\sum_{ij}\pi_{ij}\ln\left(\frac{\pi_{ij}}{p_{i}}\right)$ (42)
$=-\sum_{ij}\delta_{i,j}p_{i}\ln\left(\frac{p_{i}}{p_{i}}\right)=-\sum_{i}p_{i}\ln(1)=0.$
Notice that the requirement for thermal tensor demanded above can be guaranteed not only in zero or infinite temperature but is satisfied, for example, by construction (36). Additional discussion and examples of correlating properties of LTOCC in finite temperature are presented in Appendix C.
The correlation-generating power of LTOCC increases dramatically if we allow retention of memory between rounds. Suppose we allow ourselves to use one round of parallel LTOCC, especially $\mathcal{S}$LTOCC, which is a subset of two rounds of LTOCC with memory. In that case the above observation can be extended to encompass non-sharp states, and even Gibbs state.
@@
Consider a $\mathcal{S}LTOCC$ operation consisting of two copies of the same extremal (bi)thermal tensor $M_{ij,kl}=T_{ikl}T_{jkl}$, which entries are either 0 or $1$. Then for any product input state $\mathbf{p}\otimes\mathbf{q}$ the output state $\mathbf{\pi}=M(\mathbf{p}\otimes\mathbf{q})$ is maximally correlated.
@@
Let us denote $T(\mathbf{p},\mathbf{q})=\mathbf{r}$, then, since in each column ($\forall_{j,k}$) of $T_{ijk}$ there is only a single entry equal to $1$ and the rest of them are 0, we get
$\pi_{ij}$ $=\sum_{kl}M_{ij,lk}p_{k}q_{l}=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{kl}\delta_{i,j}T_{ijk}p_{k}q_{l}$ (43)
$=\delta_{ij}r_{i}\penalty 10000\ .$
@@
VI.2 Thermally restricted Bell nonlocality
[…]
where $E(a_{i},b_{j})$ are expectation values corresponding to measuremnt settings $a_{i},b_{j}$ for Alice and Bob, respectively. The measurement outcomes are further restricted to $\pm 1$.
[…]
Observation 8.
Consider a pair of systems of local dimension $d$ with non-degenerate local Hamiltonians, capable of implementing LTOCC. In this setting it is impossible to break classical bound of inequality (44).
Proof.
The set of free measurements of LTOCC consists of incoherent measurements, which are compatible. At the same time Bell nonlocality can be detected if and only if the measurements are incompatible [73], leading to contradiction. ∎
Thus, in the single-copy energy-nondegnerate scenario it is impossible to detect breaking of any Bell nonlocality when restricted to LTOCC.
[…]
$A^{\prime}_{i}=\mathbb{I}_{1}\oplus A_{i}\oplus\mathbb{I}_{1}=\matrixquantity(1&0&0\\ 0&A_{i}&0\\ 0&0&1)$
as observables implementable by two-outcome measurement local to Alice, and similarly $B^{\prime}_{j}$ for Bob. Direct calculation shows that
[…]
for any $A^{\prime}_{i}$ and $B^{\prime}_{j}$ realisable with incoherent measurement and local thermal operations. This leads to a general theorem.
Theorem 13.
[…]
with $D_{deg}$ being the sum of dimensions of energy-degenerated subspaces and $D_{ndeg}$ sum of dimensions of non-degenerated subspaces. In particular for Hamiltonian with equal energy spacing $D_{\text{ndeg}}=2,\;D_{\text{deg}}=d^{n}-2$ and for Hamiltonian with rationally-independend energy spacing $D_{\text{ndeg}}=d,\;D_{\text{deg}}=d^{n}-d$.
Proof.
[…]
$A_{i}=\Pi_{\text{ndeg}}\oplus\quantity(\bigoplus_{k}A_{i,k})$ (48)
[…]
Corollary 4.
The same bound (47) holds for arbitrary $n$-copy entangled state $\rho^{\otimes n}$.
Proof.
[…]
Interestingly, the statement extends to local thermal operations aided by both communication and memory.
Theorem 14.
[Extension of Theorem 13] Bound (47) on the CHSH operator holds also for local thermal operations extended by communication and memory [LTOCC(+M)]
Proof.
For the clarity of the discussion the proof of this Theorem, toghether with necesery Lemmas, is presented in the Appendix E.
Theorem 14 shows three distinct critical values of the CHSH scenario:
[…]
This leads to a proposal of a new thermally-restricted CHSH (ThCHSH) scenario, in which Alice and Bob are tasked with maximising value of CHSH correlator using an $n$-copy entangled state. When no memory is involved, the central evaluating party, Charlie, can verify that Alice and Bob have access to non-thermal operations by verifying breaking of the bound (47). The scenario can even be extended by memory shared between Alice and Bob, which can be used to communicate results of intermediate measurements under LTOCC+M paradigm; inspection of memory state by Charlie can be used to verify that parties have not communicated the measurement settings. While being a proof-of-principle rather than loophole-free robust scenario, it provides a way to certify use of resources via Bell-like experiments.
Figure 9 demonstrates the generic properties of convergence of the value of three bounds presented in eq. (49). Although specific values depend on the strucutre of energy gaps, showing higher violation when the spectrum is $\mathbb{Q}$-dependent, the general convergence behaviour is dependent primarily on the local dimensions, with behaviour of convergence to Tsirelson bound as $\mathcal{O}(d^{-n})$, with multiplicative constant depending explicitly on the number of energy-nondegenerate subspaces.
Finally, let us recall that we focused on CHSH mainly for the sake of concreteness. The arguments corresponding to energy-degenerate subspaces are, however, fully general, and thus we put forward the following general statement
Theorem 15.
Consider an arbitrary bipartite Bell scenario with an expectation value $S$ bound by classical and quantum bounds $\mathcal{B}_{\text{cl}}$ and $\mathcal{B}_{\text{q}}$. A realisation using $n$ copies of $d$-dimensional state under the same restrictions as in Theorem 13
$S\leq\frac{\mathcal{B}_{\text{cl}}D_{\text{ndeg}}+\mathcal{B}_{\text{q}}D_{\text{deg}}}{d^{n}}.$ (50)
VII Summary and Outlook